OpenStudy (anonymous):

State the slope and the y-intercept for the graph of the equation. x + y = 5 a. 5;-1 c. -1;6 b. -1; 5 d. -1; 4

OpenStudy (butterflydreamer):

you have your equation " x + y = 5" We want to change this into slope-intercept form first :) Slope intercept form is y = mx + b So x + y = 5 , to change this, simply subtract x from both sides.

OpenStudy (anonymous):

thank you~ i'm having a major brain fart right now

OpenStudy (anonymous):

so that would be 5x+y.. ?

OpenStudy (anonymous):

and the y= into -1? 5;-1

OpenStudy (butterflydreamer):